Find the slope, if it exists, of the line.
please explain and show me how to do these problems, ty
y= -10 + 7
y=10/3X -5/7

Answer by mananth(16949)

y=-10 +7
y=-3
..
this is a line parallel to the x axis. Hence no slope.
...
y=(10/3 )x -5/7
compare this equation with the equation
y= mx+b
compare the co-efficients of x
m = (10/3)